OpenStudy (anonymous):

how to find inflection point of f(x)=4x^2 +2x +1 ?

OpenStudy (marissalovescats):

You will have to find the 2nd derivative, set it = to 0 and solve for x

OpenStudy (marissalovescats):

Or graph it and see where it crosses the x axis at.

OpenStudy (anonymous):

the second derivative is 8 though.

OpenStudy (anonymous):

so im guessing there is no inflection point ? o-o

OpenStudy (imstuck):

It's a parabola so it does have an inflection point, but the graph does not go through the x axis, so the solutions are imaginary. The graph has a min value at the inflection point.

OpenStudy (anonymous):

is there a relative minimum at -1/4 ?

OpenStudy (imstuck):

The min of the graph occurs at (0, 1).

OpenStudy (imstuck):

That is where the graph reaches its lowest point and then begins to go back up.

OpenStudy (anonymous):

no.

OpenStudy (imstuck):

No what?

OpenStudy (anonymous):

i just graphed it.

OpenStudy (imstuck):

So did I.

OpenStudy (anonymous):

oh Lol

OpenStudy (anonymous):

well use this it will help you http://www.wolframalpha.com/input/?i=4x%5E2+%2B2x+%2B1&dataset=

jimthompson5910 (jim_thompson5910):

IMStuck, all parabolas are either concave up or concave down throughout their entire domain. Since you have no change in concavity, you won't have an inflection point on a parabola.

OpenStudy (anonymous):

@IMStuck the graph reaches its lowest point at -1/4, 3/4 not 0,1 you graphed it wrong ?

OpenStudy (anonymous):

@jim_thompson5910 thank you... so it has no inflection point right..?

jimthompson5910 (jim_thompson5910):

correct here's how you show it through derivatives f(x) = 4x^2 + 2x + 1 f ' (x) = 8x + 2 f '' (x) = 8 The second derivative is always positive. So f(x) is always concave up. There are no changes in concavity (ie f ''(x) doesn't change in sign at all), so there are no inflection points.

OpenStudy (anonymous):

@jim_thompson5910 thanks for clarifying, also there is a relative minimum at -1/4 ? o-o

jimthompson5910 (jim_thompson5910):

f ' (x) = 8x + 2 0 = 8x + 2 -2 = 8x x = -2/8 x = -1/4 This is a critical value. Because f ''(x) > 0 for all x, this means f(x) is concave up which leads to a relative min (turns out it's actually an absolute min too)

jimthompson5910 (jim_thompson5910):

So yes, you are correct

OpenStudy (anonymous):

Okay thanks, that other guy confused me I thought I got it wrong!

jimthompson5910 (jim_thompson5910):

you're welcome